An insurance company insures a person's antique coin collection worth $20,000 for an annual premium of $300. If the company figures that the probability of the collection being stolen is 0.002, what will be the company's expected profit?

PLEASE CREATE CHART
Answer #1

The chart is here -

X P(X)
300 0.998
0 0.002

E(X) = Σ X.P(X) = 300*0.998 + 0*0.002 = 299.4 + 0 = 299.4

The expected profit is $299.4.